mary wants to make 20 gal. of a 52% acid solution by mixing together a 84% acid solution and a 4% acid solution. how much of each must she use?

Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39617)

Only if density does not matter, you can simply follow this example:
single variable for two-part mixture - video example

Answer by Alan3354(69443)

mary wants to make 20 gal. of a 52% acid solution by mixing together a 84% acid solution and a 4% acid solution. how much of each must she use?
----------------------
e = amount of 84%
f = amount of 4%
-----
e + f = 20 (total solution)
84e + 4f = 20*52 (total acid)
-----